OpenStudy (anonymous):

Just for verification, can you please check my answer? Question: 3(x+2) = 6 ------ 2(x+1) my answer: 6(x^2+3x+1) Work: expanded first after multiplying both sides with 2(x+1).==> 6x^2+18+6==> took 6 out. to get final answer. Am i correct

OpenStudy (anonymous):

What are your instructions?

OpenStudy (anonymous):

Solve for x... shoot that means its incomplete.

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

im stuck :( whats the next step? i cant think of any factors of 6x^2+18x+6... ah am i not suppose to factor, and use quadratic formula?

OpenStudy (anonymous):

Hang on, what is the equation you just posted equal to?

OpenStudy (anonymous):

my previous post is the multiplication of both sides with 2x+2

OpenStudy (anonymous):

You have 6x^2+18x+6=0 right now, right?

OpenStudy (anonymous):

like (3x+6)(2x+2) = 6 =>6x^2+18+12 = 6 ==> 6x^2+18+6

OpenStudy (anonymous):

yes = 0

OpenStudy (anonymous):

Perfect, so then you can factor out the 6 like you did before and use the quadratic formula on what's in the parentheses

OpenStudy (anonymous):

ok one sec.

OpenStudy (anonymous):

ok so it would be ___ ___ 3 ±√-1 and 3 ±√-1 ------ ------- 2 2

OpenStudy (anonymous):

just the quadratic formula answers

OpenStudy (anonymous):

*First one = + not ± *Second one = - not ±

OpenStudy (anonymous):

Can you show me how you set up the quadratic formula because I got \[(-3\pm \sqrt{5})/2\] for the answer

OpenStudy (anonymous):

umm ok -b ± √-b - 4ac --------------- 2a so -3±√3-4(1*1) / 2(1) * hey im sorry but can you help me in an hour? im at work and i over procrastinated and i need to finish my work, but ill be back online within an hour or hour and a half.

OpenStudy (anonymous):

I gotta go soon, but FYI your formula's wrong: it's supposed to be b^2-4ac under the radical

OpenStudy (anonymous):

Oh im sorry that was a typo...in both.

OpenStudy (anonymous):

Ah i see, i did this within my previous answer too.

OpenStudy (anonymous):

Does that solve the problem?

OpenStudy (anonymous):

yes, but what would my final answer look like? 6 ((-3±√-5)/2) ?

OpenStudy (anonymous):

No, because the 6 we factored out earlier doesn't affect the final answer because if we divided both sides by 6 before using the quadratic formula, we'd get x^2+3x+2=0 since 0/6=0

OpenStudy (anonymous):

I hope that made some sense

OpenStudy (anonymous):

So, your final answer would be \[(-3\pm \sqrt{-5})/2\]

OpenStudy (anonymous):

Because the 6 canceled earlier

OpenStudy (anonymous):

hmm ok i think i understand. because we divided 6, theres no need for its return. so i would just write x = (-3±√-5)/2

OpenStudy (anonymous):

Yes! :)

OpenStudy (anonymous):

Unless your book/teacher want you to write the two answers separately

OpenStudy (anonymous):

In which case it would be \[x=(-3+ \sqrt{-5})/2\] and \[x=(-3-\sqrt{-5})/2\]

OpenStudy (anonymous):

Any more questions?

OpenStudy (anonymous):

Umm that's all for now, i should do my work now. But thank you very much :) I'll most likely post more questions within an hour or so to check my work so make sure i don't make mistakes like this again :p

OpenStudy (anonymous):

Thanks again, bye

OpenStudy (anonymous):

No prob. Bye